The longest side of a right triangle is 3 ft longer than three times the length of the shortest side. The other side of the triangle is 3 ft shorter than three times the length of the shortest side.
Find the lengths of the shortest, longest and the other side of the triangle.

Answer by Boreal(15235)

short side is x
longest side, the hypotenuse, is 3x+3
the other side is 3x-3
Therefore, x^2+(3x-3)^2=(3x+3)^2.
x^2+9x^2-18x+9=9x^2+18x+9
x^2-36x=0
x(x-36)=0
x=36 ft, the shortest side
the other side is 105 ft.
the hypotenuse is 111 ft.
divisible by 3, it is a 12-35-37 triangle
12^2=144
35^2=1225
37^2=1369, which is the sum of the first two.